# Contraction of Private Quantum Channels and Private Quantum Hypothesis Testing

### v1

@@ IV-A Quantum Hypothesis Testing with No Privacy Constraints
$\max\!\left\{\frac{\ln(pq/\alpha)}{-\ln F(\rho,\sigma)},\frac{pq-\alpha(1-\alpha)}{pq\left[d_{\mathrm{B}}(\rho,\sigma)\right]^{2}}\right\}\leq\mathrm{SC}_{(\rho,\sigma)}(\alpha,p,q)\leq\left\lceil\frac{2\ln\!\left(\frac{\sqrt{pq}}{\alpha}\right)}{-\ln F(\rho,\sigma)}\right\rceil.$ (114)
@@ IV-B Private Quantum Hypothesis Testing
$C_{\varepsilon,p,q,\alpha}\coloneqq\max\!\left\{\frac{\ln(pq/\alpha)(e^{\varepsilon}+1)}{\varepsilon(e^{\varepsilon}-1)},\frac{\left(pq-\alpha(1-\alpha)\right)(e^{\varepsilon}+1)}{2pq\ (e^{\varepsilon/2}-1)^{2}}\right\}.$ (119)
@@
Lower bound: To obtain the first part of the lower bound, observe that with the data processing of the channel $\mathcal{A}^{\otimes n}$ and applying [37, Proposition 5.2] we have
@@
$\mathrm{SC}^{\varepsilon}_{\rho,\sigma)}$ $\geq\inf_{\mathcal{A}\in\mathcal{B}^{\varepsilon}}\frac{\ln(pq/\alpha)}{-\ln F\!\left(\mathcal{A}(\rho),\mathcal{A}(\sigma)\right)}$ (130)
@@
$\mathrm{SC}^{\varepsilon}_{\rho,\sigma)}$ $\geq\inf_{\mathcal{A}\in\mathcal{B}^{\varepsilon}}\frac{pq-\alpha(1-\alpha)}{pq\left[d_{\mathrm{B}}(\mathcal{A}(\rho),\mathcal{A}(\sigma))\right]^{2}}$ (133)
@@
For orthogonal states, we need only one sample of the unknown state to declare whether it is $\rho$ or $\sigma$ when we have access to non-privatized samples of $\rho$ or $\sigma$. However, to achieve $\varepsilon$-QLDP, the input states are privatized by applying the channel $\mathcal{A}$, so that we have access to $\mathcal{A}(\rho)$ and $\mathcal{A}(\sigma)$ only. To this end, we need samples on the order of $\Theta\!\left((e^{\varepsilon}+1)/(e^{\varepsilon}-1)\right)$ for all $\varepsilon>0$, which is apparent from Corollary 2.
@@
Remark 6 discusses about the cost of privacy when we have access to privatized data samples instead of original data. Next, we show that when certain conditions are met, the impact of privacy on the sample complexity is not significant and is comparable to the non-private sample complexity.
@@
See Section -B. ∎
Note that from Proposition 6, when $\rho$ and $\sigma$ are qubits, the order optimality in sample complexity is achieved by the privatization mechanism presented in Proposition 1 by choosing the measurement channel to correspond the measurement on to the eigenbasis of $\rho\#\sigma^{-1}$.
@@
The proof is presented in Appendix -C. [^1] Note that this lower bound also holds in the case where we consider the same private channel for every copy of the state.
@@
Above, we considered $\mathcal{A}_{i}$ chosen independent of $\mathcal{A}_{j}$, where $i\neq j$. It would be an interesting future work to understand whether adaptive strategies could decrease the sample complexity or number of copies of the privatized states required to minimize the cost of privacy.
@@
where $\widetilde{D}_{\beta}$ is the sandwiched Rényi relative entropy of order $\beta$ [46, 47].
@@ IV-C Instance-Specific Private Channels and Sample Complexity
Furthermore, for high privacy regime $\varepsilon<1$ choosing $p,q$ to be constants, we conclude that
@@ V Contraction under $(\varepsilon,\delta)$-QLDP channels
Next, we quantify the contraction of the normalized trace distance by instantiating the quantum generalized divergence to be the normalized trace distance. The following extends Theorem 2 and provides an alternative proof for the converse bound in the proof of Theorem 2.
@@
This follows by the use of the mechanism introduced in the achievability part of the proof of Theorem 5 together with the similar techniques used in the proof of the upper bound in Theorem 4.
We leave the characterization of sample complexity of hypothesis testing when privacy is imposed by $(\varepsilon,\delta)$-QLDP for future work. In Section -E, we provide some tools that may be of relevance to accomplish this goal. In particular, we prove several connections between $\varepsilon$-QLDP and $(\varepsilon,\delta)$-QLDP mechanisms, which may potentially be used to infer about the impact of $(\varepsilon,\delta)$-QLDP constraints using the formal guarantees derived for $\varepsilon$-QLDP in Theorem 4.
@@ VI-A Quantum Fairness through QLDP
Classical decision models including machine learning models are prone to discriminating against individuals based on different characteristics, for example, skin color or gender [48]. This has even led to legal mandates of ensuring fairness. With the introduction of quantum machine learning models, there is also a risk of whether fairness will be ensured to both classical and quantum data fed into these algorithms. In [49], it is shown that noisy quantum algorithms can improve fairness. Since fairness and privacy are related domains, it is vital to understand the impact of private algorithms on ensuring fairness as well.
Formally, quantum fairness aims to treat all input states equally, meaning that all pairs of input states that are close in some distance metric (e.g., close in normalized trace distance) should yield similar outcomes when processed by a quantum channel [49]. Define $\mathcal{A}\coloneqq\mathcal{M}\circ\mathcal{E}$, which is a quantum-to-classical channel where a quantum channel $\mathcal{E}$ is followed by measurement channel comprised of a POVM $\{M_{i}\}_{i\in\mathcal{O}}$. With that, quantum fairness is defined in [49] as follows.
@@
By choosing $\hat{D}$ to be the normalized trace distance and $d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\!\left|\mathrm{Tr}\!\left[M_{i}\mathcal{E}(\rho-\sigma)\right]\right|$, [14, Proposition 14] shows that $\mathcal{A}$ satisfying $\varepsilon$-QLDP implies that it also $(\alpha,\sqrt{\varepsilon^{\prime}/2})$-fair, where $\varepsilon^{\prime}=\min\{\varepsilon,\varepsilon^{2}/2\}$. Note that for the case $\varepsilon>1$, the above bound could be weak. With Theorem 2, we improve the existing bound on achievable fairness through $\varepsilon$-QLDP mechanisms, which is applicable for all $\varepsilon\geq 0$, and we also extend them to cases for which $\delta\geq 0$.
@@ VI-B Stability for Quantum Learning through Private Channels
A learning algorithm is known to be stable if its output does not depend too much on any individual training data [50, 51]. It is also shown that stability and generalization of a classical learning algorithm to new inputs are closely related, where stability implies generalization using information-theoretic tools [51]. Classical differential private learners are proved to generalize well [51, Theorem 5]. In [5, Proposition 6], it is shown that learners satisfying a mutual-information-based variant of differential privacy also satisfy algorithmic stability, and hence generalize well for new data.
@@ VII Concluding Remarks and Future Directions
In this work, we derived upper bounds on the contraction of quantum divergences under privacy constraints imposed by channels that satisfy $\varepsilon$-QLDP. Notably, we fully characterized the privatized contraction coefficient of trace distance, proving the converse by deriving an upper bound on the contraction of the hockey-stick divergence, and proving achievability by proposing a novel QLDP mechanism that achieves the bound. Next, with the tools developed, we studied the quantum private hypothesis testing and provided bounds on the sample complexity, when we have access to privatized samples. We showcased the cost of privacy on this statistical task while providing tight sample complexity bounds on special sets of states, and special classes of private channels. In addition, we also characterized the privatized contraction coefficient of trace distance under $(\varepsilon,\delta)$-QLDP mechanisms. Finally, we explored how the contraction of divergences derived can be applied in other applications including ensuring fairness and providing formal guarantees on stability and generalization for quantum learning settings by addressing an open question posted in [35].
@@
Another interesting research question is to characterize the sample complexity of asymmetric binary quantum hypothesis testing. Here one could make use of the non-private sample-complexity bounds presented in [32, Theorem 9]. Preliminary results in this direction are found in Remark 9.
@@ Acknowledgements
We are grateful to Hao-Chung Cheng, Christoph Hirche, and Cambyse Rouzé for coordinating the arXiv post of their paper titled ”Sample Complexity of Locally Differentially Private Quantum Hypothesis Testing”. We would like to thank Behnoosh Zamanlooy and Shahab Asoodeh for sharing the full version with appendices of their conference paper [43], which inspired some of our proofs. TN acknowledges support from the NSF under grant no. 2329662. MMW acknowledges support from the NSF under grants 2329662, 2315398, 2304816.

### v2

@@ IV-A Quantum Hypothesis Testing with No Privacy Constraints
$\max\!\left\{\frac{\ln(pq/\alpha)}{-\ln F(\rho,\sigma)},\frac{1-\frac{\alpha(1-\alpha)}{pq}}{\left[d_{\mathrm{B}}(\rho,\sigma)\right]^{2}}\right\}\leq\mathrm{SC}_{(\rho,\sigma)}(\alpha,p,q)\leq\left\lceil\frac{2\ln\!\left(\frac{\sqrt{pq}}{\alpha}\right)}{-\ln F(\rho,\sigma)}\right\rceil.$ (118)
@@ IV-B Private Quantum Hypothesis Testing
$C_{\varepsilon,p,q,\alpha}\coloneqq\max\!\left\{\frac{\ln(pq/\alpha)(e^{\varepsilon}+1)}{\varepsilon(e^{\varepsilon}-1)},\frac{\left({1-\frac{\alpha(1-\alpha)}{pq}}\right)(e^{\varepsilon}+1)}{2\ (e^{\varepsilon/2}-1)^{2}}\right\}.$ (123)
@@
Lower bound: To obtain the first part of the lower bound, observe that with the data processing of the channel $\mathcal{A}^{\otimes n}$ and applying [38, Proposition 5.2] we have
@@
$\mathrm{SC}^{\varepsilon}_{{(}\rho,\sigma)}$ $\geq\inf_{\mathcal{A}\in\mathcal{B}^{\varepsilon}}\frac{\ln(pq/\alpha)}{-\ln F\!\left(\mathcal{A}(\rho),\mathcal{A}(\sigma)\right)}$ (134)
@@
$\mathrm{SC}^{\varepsilon}_{{(}\rho,\sigma)}$ $\geq\inf_{\mathcal{A}\in\mathcal{B}^{\varepsilon}}\frac{pq-\alpha(1-\alpha)}{pq\left[d_{\mathrm{B}}(\mathcal{A}(\rho),\mathcal{A}(\sigma))\right]^{2}}$ (137)
@@
For orthogonal states, we need only one sample of the unknown state to declare whether it is $\rho$ or $\sigma$ when we have access to non-privatized samples of $\rho$ or $\sigma$. However, to achieve $\varepsilon$-QLDP, the input states are privatized by applying the channel $\mathcal{A}$, so that we have access to $\mathcal{A}(\rho)$ and $\mathcal{A}(\sigma)$ only. To this end, we need samples on the order of $\Theta\!\left(\left((e^{\varepsilon}+1)/(e^{\varepsilon}-1)\right)^{2}\right)$ for all $\varepsilon>0$, which is apparent from Corollary 2.
@@
Remark 6 discusses the cost of privacy when we have access to privatized data samples instead of original data. Next, we show that when certain conditions are met, the impact of privacy on the sample complexity is not significant and is comparable to the non-private sample complexity. This also provides a sense of how large the privacy parameter $\varepsilon$ should be in order to achieve a similar sample complexity as in the non-private case, where we do not need to pay an extra cost to ensure privacy, $d_{B}(\rho,\sigma)>1$.
@@
See Section -D. ∎
Note that from Proposition 7, when $\rho$ and $\sigma$ are qubits, the order optimality in sample complexity is achieved by the privatization mechanism presented in Proposition 1 by choosing the measurement channel to correspond to the measurement in the eigenbasis of $\rho\#\sigma^{-1}$.
@@
The proof is presented in Appendix -E. [^1] Note that this lower bound also holds in the case where we consider the same private channel for every copy of the state.
@@
Above, we considered $\mathcal{A}_{i}$ chosen independently of $\mathcal{A}_{j}$, where $i\neq j$. It is an interesting question for future work to understand whether adaptive strategies can decrease the sample complexity or number of copies of the privatized states required to minimize the cost of privacy.
@@
where $\widetilde{D}_{\beta}(\omega\|\tau)\coloneqq\frac{1}{\beta-1}\ln\operatorname{Tr}[(\tau^{(1-\beta)/2\beta}\omega\tau^{(1-\beta)/2\beta})^{\beta}]$ is the sandwiched Rényi relative entropy of order $\beta$ [49, 50].
Remark 10 (Comparison to [36]).
The key difference of our findings and the independent work [36] is that our sample complexity bounds in Theorem 4 hold for arbitrary $\alpha,p,q$ such that $\alpha\in(0,pq)$, while their results are tailored for specific values (see Theorem 4.10 and Eq. (1.6) and Eq. (1.7) in [36]). If we also choose those specific values that they consider, our upper bound on the sample complexity coincides with theirs. For the lower bound, our bound is tighter when considering the dependence on the privacy parameter $\varepsilon$ while their bound is tighter when considering the dependence on $T(\rho,\sigma)$. Our lower bound is a key tool in precisely characterizing the sample complexity of orthogonal quantum states when privatized quantum states are given to us, as we have concluded in Corollary 2.
@@ IV-C Instance-Specific Private Channels and Sample Complexity
Furthermore, for the high privacy regime $\varepsilon<1$ and choosing $p,q$ to be constants, we conclude that
@@
In summary, this section shows that there exist sets of private channels that depend on the states chosen (defined in (163)), where the general sample complexity bounds stated in Theorem 4 can be improved as in Proposition 8 to minimize the gap between the lower and upper bounds. We leave further analysis on practically relevant and operationally motivated sets of channels achieving similar improvements for future work.
@@ V Contraction under $(\varepsilon,\delta)$-QLDP channels
Next, we quantify the contraction of the normalized trace distance by instantiating the generalized divergence to be the normalized trace distance. The following extends Theorem 2 and provides an alternative proof for the converse bound in the proof of Theorem 2.
@@
Note that Theorem 5 also improves the best known bound for the contraction of total variation distance in the classical setting (See [51, Lemma 1]).
@@
This follows by the use of the mechanism introduced in the achievability part of the proof of Theorem 5 together with similar techniques used in the proof of the upper bound in Theorem 4. For the setting of instance specific private channels in Section IV-C, it is possible to extend Proposition 8 to the $\delta\neq 0$ setting with the use of the contraction of trace distance in Theorem 5.
We leave the general characterization of sample complexity of hypothesis testing when privacy is imposed by $(\varepsilon,\delta)$-QLDP for future work. The main technical tools needed to obtain the lower bounds as given in Theorem 4 are the characterization of the contraction of quantum relative entropy and Bures distance under the privacy constraints imposed by $(\varepsilon,\delta)$-QLDP. In Section -G, we provide some tools that may be of relevance to accomplish this goal. In particular, we prove several connections between $\varepsilon$-QLDP and $(\varepsilon,\delta)$-QLDP mechanisms, which may potentially be used to infer about the impact of $(\varepsilon,\delta)$-QLDP constraints using the formal guarantees derived for $\varepsilon$-QLDP in Theorem 4.
@@ VI-A Quantum Fairness through QLDP
Classical decision models, including machine learning models, are prone to discriminating against individuals based on different characteristics, for example, skin color or gender [52]. This has even led to legal mandates of ensuring fairness. With the introduction of quantum machine learning models, there is also a risk of whether fairness will be ensured to both classical and quantum data fed into these algorithms. In [53], it was shown that noisy quantum algorithms can improve fairness. Since fairness and privacy are related domains, it is vital to understand the impact of private algorithms on ensuring fairness as well.
Formally, quantum fairness aims to treat all input states equally, meaning that all pairs of input states that are close in some distance metric (e.g., close in normalized trace distance) should yield similar outcomes when processed by a quantum channel [53]. Define $\mathcal{A}\coloneqq\mathcal{M}\circ\mathcal{E}$, which is a quantum-to-classical channel where a quantum channel $\mathcal{E}$ is followed by a measurement channel comprised of a POVM $\{M_{i}\}_{i\in\mathcal{O}}$. With that, quantum fairness is defined in [53] as follows.
@@
By choosing $\hat{D}$ to be the normalized trace distance and $d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\!\left|\mathrm{Tr}\!\left[M_{i}\mathcal{E}(\rho-\sigma)\right]\right|$, [14, Proposition 14] states that $\mathcal{A}$ satisfying $\varepsilon$-QLDP implies that it is also $(\alpha,\sqrt{\varepsilon^{\prime}/2})$-fair, where $\varepsilon^{\prime}=\min\{\varepsilon,\varepsilon^{2}/2\}$. Note that for the case $\varepsilon>1$, the above bound could be weak. With Theorem 2, we improve the existing bound on achievable fairness through $\varepsilon$-QLDP mechanisms, which is applicable for all $\varepsilon\geq 0$, and we also extend them to cases for which $\delta\geq 0$.
@@ VI-B Stability for Quantum Learning through Private Channels
Designing learning algorithms that also ensure privacy for input data is of importance, and it has been widely studied in the classical setting (for example, see [54]). A learning algorithm is known to be stable if its output does not depend too much on any individual training data [55, 56]. It is also known that stability and generalization of a classical learning algorithm to new inputs are closely related, where stability implies generalization using information-theoretic tools [56]. Classical differentially private learners were proved to generalize well [56, Theorem 5]. In [5, Proposition 6], it was shown that learners satisfying a mutual-information-based variant of differential privacy also satisfy algorithmic stability, and hence generalize well for new data.
@@ VII Concluding Remarks and Future Directions
In this work, we derived upper bounds on the contraction of quantum divergences under privacy constraints imposed by channels that satisfy $\varepsilon$-QLDP. Notably, we fully characterized the privatized contraction coefficient of trace distance, proving the converse by deriving an upper bound on the contraction of the hockey-stick divergence, and proving achievability by proposing a novel QLDP mechanism that achieves the bound. Next, with the tools developed, we studied quantum private hypothesis testing and provided bounds on the sample complexity, when we have access to privatized samples. We showcased the cost of privacy on this statistical task while providing tight sample complexity bounds on special sets of states, and special classes of private channels. In addition, we also characterized the privatized contraction coefficient of trace distance under $(\varepsilon,\delta)$-QLDP mechanisms. Finally, we explored how the contraction of divergences derived can be applied in other applications including ensuring fairness and providing formal guarantees on stability and generalization for quantum learning settings by addressing an open question posed in [35].
@@
Another interesting research question is to characterize the sample complexity of asymmetric binary quantum hypothesis testing. Here one could make use of the non-private sample-complexity bounds presented in [32, Theorem 9]. Preliminary results in this direction are found in Remark 9. It is also an interesting future research direction to understand the impact of general noisy channels on the sample complexity of quantum hypothesis testing. Towards achieving this, main technical tool needed is the contraction coefficient of quantum divergences under general noisy channels.
@@ Acknowledgements
We thank Kaiyuan Ji for helpful discussions. We are grateful to Hao-Chung Cheng, Christoph Hirche, and Cambyse Rouzé for coordinating the arXiv post of their paper titled ”Sample Complexity of Locally Differentially Private Quantum Hypothesis Testing”. We would like to thank Behnoosh Zamanlooy and Shahab Asoodeh for sharing the full version with appendices of their conference paper [46], which inspired some of our proofs. TN acknowledges support from the NSF under grant no. 2329662. MMW acknowledges support from the NSF under grants 2329662, 2315398, 2304816.